OpenStudy (anonymous):

What is the equation of the line that passes through (4, -1) and (-2, 3)? 2x - 3y - 5 = 0 2x + 3y - 5 = 0 -2x + 3y - 5 = 0

OpenStudy (anonymous):

Use this : m=Y2-Y1/X2-X1 X1=4 Y1=-1 We just want (-2,3) Caz formula is this : Y2-Y=m(X2-X) X2=-2 Y2=3 Now answer ! (X and Y are not X1 and Y1 ) Got it ?:)

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

You re welcome ! Now look at the answer : m=-1/3 and : 3-y=1/3(-2-x) 3-y=-2/3-1/3x y=3+2/3+1/3x y=11/3+1/3x

OpenStudy (anonymous):

But in the answers we dont have this ! You should find the same answer !:)